OpenStudy (anonymous):

WILL BEST ANSWER A number is decreased by seven. Then, the new number is multiplied by 3 to get an answer of -9. What is the original number? -30 30 4 -4 So confusing? How do I work it out?

OpenStudy (abdullahm):

So to work it out, we need to work backwords. They said our final answer was -9. And we need to find the original number. So we have to go backwords with the steps doing the opposite of what they had originally did to get to the first number. Do you understand?

OpenStudy (anonymous):

Yeah a little, how would I work it backwards?

OpenStudy (abdullahm):

So for example, A number is multiplied by 4, and now we get 20 So to work this backwords, we have to divide 20 by 4 to know which number when multiplied by 4 got us to 20. And that is 5. That is just an example. Does that make it more clear to you? (:

OpenStudy (anonymous):

Yes, Thank you!

OpenStudy (abdullahm):

So, the question is: A number is decreased by seven. Then, the new number is multiplied by 3 to get an answer of -9. So we have to work one step at a time backwords until we reach the beginning of the question. So what number multiplied by 3 would get us -9?

OpenStudy (anonymous):

Thinking...

OpenStudy (anonymous):

we would multiply by 3 on the answer choices to see which one equals -9 right? @AbdullahM

OpenStudy (abdullahm):

not in the answer choices, just divide -9 by 3 to find out what the previous number was. We're working backwords and once we finish all the steps we will get the answer which is one of the answer choices.

OpenStudy (anonymous):

so... We divide the answer choices by -9 and then multiply by 3?

OpenStudy (anonymous):

@AbdullahM

OpenStudy (abdullahm):

we have nothing to do with the answer choices right now.

OpenStudy (abdullahm):

So, first tell me what is \(\sf\Large \frac{-9}{3}=?\)

OpenStudy (anonymous):

@AbdullahM -3

OpenStudy (abdullahm):

Ok, now we move back to the step before it which is: A number is decreased by seven

OpenStudy (abdullahm):

So lets say the original number is 'n'. Then that means n - 7 = -3

OpenStudy (anonymous):

4!!!!!

OpenStudy (anonymous):

Thank you! @AbdullahM

OpenStudy (abdullahm):

Good job! :D

OpenStudy (anonymous):

:)

OpenStudy (abdullahm):

And just ot make sure, you can use 4 and work through the steps like normal to make sure we arrive at -9

OpenStudy (anonymous):

Thank you so much!

OpenStudy (abdullahm):

A number is decreased by seven. Then, the new number is multiplied by 3 to get an answer of -9. So we have 4 - 7 = -3 and then -3 x 3 = -9 So it works, and is your correct answer :)

OpenStudy (anonymous):

Okay Thank you!